8. Consider the following data: Autonomous Consumption 600 600 600 Income 10000 11000 12000 Taxation 2000 2000 2000 Transfers Consumption 400 400 400 6240 6940 7640 a) What is the marginal propensity to consume? -0.1 △4 /LO00-10000-l, b) What is the value of Consumption for an Income equal 11,500? 0O7 5OO 400- 2000) (-7,ら3D Pencil marks are my attempt. Please ignore.
Answer #1

Answer: a) We know that consumption is defined as C= \bar C + c(Y_d) \ where \ Y_d = Y - Tax + Transfer

for marginal propencity to consume we will diffrentiate the consumption function with respect to Yd.

MPC= MPC= \Delta C/ \Delta Y_d

Using any two period consumption we know that

MPC= (6940-6240)/ (11000-10000) = 700/1000= 0.7

b) Using the above formula we can calculate the consumption as

C= 600+ 0.7*(11500-2000+400) = 7530

So your answer are absolutely corrects.